# InquireMobile: Teaching VLM-based Mobile Agent to Request Human Assistance via Reinforcement Fine-Tuning

### v1

@@ Abstract
Recent advances in Vision-Language Models (VLMs) have enabled mobile agents to perceive and interact with real-world mobile environments based on human instructions. However, the current fully autonomous paradigm poses potential safety risks when model understanding or reasoning capabilities are insufficient. To address this challenge, we first introduce InquireBench, a comprehensive benchmark specifically designed to evaluate mobile agents’ capabilities in safe interaction and proactive inquiry with users, encompassing 5 categories and 22 sub-categories, where most existing VLM-based agents demonstrate near-zero performance. In this paper, we aim to develop an interactive system that actively seeks human confirmation at critical decision points. To achieve this, we propose InquireMobile, a novel model inspired by reinforcement learning, featuring a two-stage training strategy and an interactive pre-action reasoning mechanism. Finally, our model achieves an 46.8% improvement in inquiry success rate and the best overall success rate among existing baselines on InquireBench. We will open-source all datasets, models, and evaluation codes to facilitate development in both academia and industry.
{strip}[Uncaptioned image]
@@
1 Introduction
Recent advances in Vision-Language Models (VLMs) have significantly propelled the capabilities of mobile agents, enabling them to autonomously perceive and interact with complicated real-world environments based on human instructions (Hurst et al. 2024; Bai et al. 2025; Li and Huang 2025). Traditionally, these mobile agents (Hong et al. 2024; Qin et al. 2025) adopt a fully autonomous paradigm: after receiving user instructions, the mobile agent independently interacts with its mobile environment to accomplish designated tasks.
However, this paradigm presumes “absolute trust” in the agent’s decision-making abilities—a presupposition that becomes problematic when model understanding or reasoning capacity is insufficient. In such cases, unverified autonomy can result in severe consequences, especially in high-stakes scenarios such as online payments, order placements, or the handling of sensitive personal data as shown in Figure 1. For instance, a misinterpreted instruction at a payment interface could lead to unintended financial transactions.
@@ 1 Introduction
Name Eval Mode # Tasks # Apps Language Inquire Data Online
AITW (Rawles et al. 2023) static - - EN
AndroidControl (Li et al. 2024) static - - EN
AMEX (Chai et al. 2024) static - - EN
A3 (Chai et al. 2025) dynamic 201 21 EN
AppAgent (Zhang et al. 2025a) dynamic 45 9 EN
AndroidArena (Xing et al. 2024) dynamic 221 14 EN
AndroidWorld (Rawles et al. 2024) dynamic 116 20 EN
Android-Lab (Xu et al. 2024) dynamic 138 9 EN
Mobile-Env (Zhang et al. 2023) dynamic 224 15 EN
CAGUI (Zhang et al. 2025b) static 603 - CN
InquireBench (Ours) dynamic 173 37 EN&CN
To bridge this research gap, we introduce InquireBench, a novel benchmark specifically designed to evaluate a mobile agent’s ability to inquire and interact safely with users. Specifically, we systematically construct diverse and challenging scenarios by leveraging both existing instructions and a vast new set of instructions generated by GPT-4o. By simulating “random walks” on real mobile phones, we trigger a breadth of potential inquire cases across 5 categories (i.e., intent confirmation, privacy and security, risk scenarios, combination and others), for evaluating user-agent interactive performance. Surprisingly, comprehensive evaluation of both open-source and closed-source VLM-based agents (Bai et al. 2025; Luo et al. 2025; Lu et al. 2025a) on InquireBench reveals a near-zero baseline performance, underscoring the urgent need for new techniques tailored for user-agent collaborative situations.
To achieve this goal, we propose InquireMobile, a model designed to teach VLM-based mobile agents to request human assistance through reinforcement fine-tuning. This new mobile agent model employs a two-stage training strategy: it begins with supervised fine-tuning (SFT) for robust format acquisition, and is followed by Group Relative Policy Optimization (GRPO) (Shao et al. 2024) training to enhance the model’s reasoning and thinking capabilities. Crucially, InquireMobile incorporates an interactive pre-action reasoning mechanism, where the agent proactively inquires from the user before executing critical actions—enabling informed decision-making through human-in-the-loop interaction. In our experiments, InquireMobile achieves a significant performance gain of 46.8% points in inquiry success rate and the best success rate over existing baselines on InquireBench, demonstrating the feasibility and necessity of proactive user engagement in agent-driven automation.
@@ 2.1 VLM-based GUI Agents
With advances in computer vision and vision-language models (VLMs), agents can directly utilize screen-visual information—such as screenshots—to better perceive and interact with the on-screen environment. Recent studies have shifted from complex framework designs to exploring end-to-end agent strategies based on VLMs, enhancing the perception and planning capabilities of VLMs. For example, CogAgent (Hong et al. 2024) presents a vision-language foundation model specializing in GUI understanding and planning, relying solely on visual inputs. UI-TARS (Qin et al. 2025) continually enhances Qwen-2-VL (Wang et al. 2024) in perception, reasoning, and memory by training on approximately 50 billion tokens. OS-Kairos (Cheng et al. 2025) employs a confidence-driven strategy to seek help from advanced models or humans when actions are unreliable, ensuring smooth task progress.
@@ 2.2 RFT-tuned GUI Agents
Inspired by the reinforcement fine-tuning (RFT) approach of DeepSeek-R1 (Guo et al. 2025), RFT-tuned VLMs (Zhou et al. 2025a; Liu et al. 2025b; Chen et al. 2025) have demonstrated excellent performance across various vision tasks. By combining SFT with reinforcement learning (e.g., GRPO), rule-based reward mechanisms significantly enhance the models’ reasoning and generalization capabilities. UI-R1 (Lu et al. 2025b) and GUI-R1 (Luo et al. 2025) extend this strategy to GUI agents, employing simple rule-based reward functions to evaluate the correctness of actions, achieving notable performance improvements with only limited datasets. InfiGUI-R1 (Liu et al. 2025a) proposes a reasoning-oriented, two-stage training paradigm, designed to gradually evolve the agent from a reactive actor to a deliberate planner. GUI-G1 (Zhou et al. 2025b) enhances the GUI interaction ability of VLMs through unified action-space rule modeling. Mobile-R1 (Gu et al. 2025) further introduces an interactive multi-round reinforcement learning framework with task-level rewards, significantly improving the agent’s exploration and error-correction abilities.
Although these models have achieved impressive performance on existing benchmarks (Cheng et al. 2024; Li et al. 2024, 2025), they focus on action execution while neglecting the interactive feedback from users. This poses significant challenges to the user privacy and security of GUI agents.
@@ 3.1.1 Triggering the Inquiry Situation
Existing datasets are typically constructed under idealized conditions and fail to capture the diverse range of anomalies encountered in real-world deployment. Traditional dataset construction methods typically rely on manual processes, including task design, execution planning, and human demonstration. However, such approaches are not suitable for our benchmark, as the interactive steps in our setting involve inherent randomness and unknown factors. Inspired by RevAct (Yang et al. 2025), we adopt a novel data collection strategy that reverses the conventional workflow and enables semi-automated dataset generation. Specifically, we first collect a large number of screenshots by simulating “random walks” on real mobile phones [^1] The devices used for data collection include Huawei and Xiaomi phones, both equipped with Android systems., which better reflects the natural interactions observed in daily life. Notably, to account for the significant differences in categories and usage patterns between Chinese and English apps, users are instructed to interact with each type of app separately. Each app is then evaluated within its respective system language setting.
After the “random walks”, we obtain approximately 80,345 screenshots. Human annotators are then asked to identify the interactive screenshots that require human intervention when agents encounter these screens, such as account login or payment execution. The interactive scenarioed are summarized into five categories:
  • Intent Confirmation: When the agent is unable to determine the next action (e.g., pop-up windows, advertisements, or ambiguous user instructions), it proactively seeks guidance from the user.
  • Privacy and Security: Actions such as login or permission granting require explicit user confirmation.
@@ 3.1.2 Thinking Generation
Annotators assess whether each item is appropriate and reasonable for the given context, and then decide whether to retain the original content, edit any unreasonable parts, or rewrite them entirely if necessary. They evaluate the model’s “thinking” for reasonableness using the format shown in Figure 4. Redundant or incorrect data is rewritten.
Figure 4: Interactive thinking format. Red denotes the user instruction, blue indicates the current state or interface, and purple shows the reason and category in inquiry scenarios.
@@ 3.1.3 Statistics
The category distribution is shown in Figure 3. More details can be found in Appendix 7.3.
@@ 4 InquireMobile
Our InquireMobile adopts a two-stage training strategy: (1) format finetuning by supervised fine-tuning (SFT), and (2) inquiry enhancement via GRPO training with verifiable rewards, to improve the agent’s interactive capabilities, as shown in Figure 5. The data used in the two training stages include both inquiry-related and general GUI human-annotated data, as described in Section 3. Given a task presented with a user instruction in natural language, the mobile agent is responsible for making action decisions to complete this instruction on the phone, if necessary, calls the user for asistance or clarification.
4.1 Preliminary
Rule-based RL can enhance the reasoning capabilities of multimodal large language models (MLLMs) through policy-based algorithms such as Group Relative Policy Optimization (GRPO) (Shao et al. 2024). By using group-normalized, token-level advantages, GRPO lowers reward sparsity and variance while removing the need for a separate value function or critic network.
In GRPO, the model generates a set of $N$ candidate responses $O=\left\{o_{1},o_{2},...,o_{N}\right\}$ for each task. Each response is evaluated by taking the corresponding actions and computing its reward $\left\{r_{1},r_{2},...,r_{N}\right\}$. Unlike PPO (Schulman et al. 2017), which relies on a single reward signal and a critic to estimate the value function, GRPO normalizes these rewards to calculate the relative advantage of each response. The relative advantage $A_{i}$ of the $i$-th response is calculated as follows:
$\begin{split}J_{\text{GRPO}}(\theta)&=\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\min\left[\frac{\pi_{\theta}(o_{i}(t)|o_{i},<t)}{\pi_{\text{old}}(o_{i}(t)|o_{i},<t)}\hat{A}_{i,t},\right.\\ &\quad\left.\text{clip}\left(\frac{\pi_{\theta}(o_{i}(t)|o_{i},<t)}{\pi_{\text{old}}(o_{i}(t)|o_{i},<t)},1-\epsilon,1+\epsilon\right)\hat{A}_{i,t}\right]\end{split}$ (2)
@@ 4.3 Stage 2: Rule-Based RL Training
Subsequently, the model was trained using GRPO with verifiable rewards. The reward function consists of format reward and action reward, defined as:
@@
Format Reward ($R_{F}$)
Following previous work (Meng et al. 2025; Gu et al. 2025; Huang et al. 2025), we introduce the format reward $R_{F}$ to encourage the model to generate structured and interpretable outputs.
@@
  • <tool_call>: The final answer in the form of a JSON object containing the function name and arguments.
@@
Action Type Reward ($R_{T}$)
$R_{T}$ assesses the the correctness of the predicted action type. In our tasks, the action space consists of nine types, including general actions such as click and swipe, as well as the inquiry action call_user. The action space is detailed in Table 5 in Appendix 7.2. $R_{T}$ is 1 if the action type exactly matches the ground truth otherwise, it is 0.
Action Argument Reward ($R_{A}$)
  • For text-based actions (e.g., type, call_user), we compute the BLEU (Papineni et al. 2002) score between the generated text and the ground truth. The BLEU score is normalized to the range $[0,1]$.
@@ 5.1 Implementation Details
The Android Studio emulator and two physical Android mobile phones serve as our primary mobile GUI interaction environments. Local monitoring scripts run on the host machine and connect to each device to manage the interaction loop.
@@
In the two training stages, we utilized 975 inquiry data to enhance interactive ability and 3,000 general GUI data to retain model’s mobile agent ability. Further more, we constructed a bilingual benchmark of 190 carefully curated instructions—95 Chinese and 95 English—each crafted to elicit uncertainty and require human-in-the-loop interaction. To mirror authentic mobile usage, each task is paired with a realistic execution environment, as detailed in Appendix 7.4.
@@
Our experiments utilize Qwen2.5-VL-3B as the base model, with GRPO implementation. The training procedure consists of two stages trained on 4 H100 GPUs. In Stage 1, we perform supervised fine-tuning (SFT) with LoRA (lora_rank =8=8) for 2 epochs with a learning rate of $1.0\times 10^{-4}$. In Stage 2, GRPO is applied for 2 epochs with 4 generations per sample and a temperature of 1 for exploration. The max steps for each task is set to 15.
@@
AppAgent (Zhang et al. 2025a), Mobile-Agent-E (Wang et al. 2025), Qwen2.5-VL-3B (Bai et al. 2025), GUI-R1-3B (Luo et al. 2025), UI-R1-3B and UI-R1-3B-E (Lu et al. 2025a) are baselines.
@@
To evaluate the performance of the existing frameworks and models in real-world scenarios, we use a Huawei phone instead of simulators as the testing environment. We evaluate the agent’s performance on these tasks using the following metrics:
  • Task Success Rate (SR): the proportion of tasks the agent completes successfully.
  • Inquiry Success Rate (ISR): the proportion of cases in which the agent makes a correct inquiry to the user at the appropriate time.
  • Task Completion Score (Score): the score of task completion rated by GPT-4o [^2] The version we used is GPT-4o-0806. Prompts provided to GPT-4o can be found in the Appendix 7.1..

### v2

@@ Abstract
Recent advances in Vision-Language Models (VLMs) have enabled mobile agents to perceive and interact with real-world mobile environments based on human instructions. However, the current fully autonomous paradigm poses potential safety risks when model understanding or reasoning capabilities are insufficient. To address this challenge, we first introduce InquireBench, a comprehensive benchmark specifically designed to evaluate mobile agents’ capabilities in safe interaction and proactive inquiry with users, encompassing 5 categories and 22 sub-categories, where most existing VLM-based agents demonstrate near-zero performance. In this paper, we aim to develop an interactive system that actively seeks human confirmation at critical decision points. To achieve this, we propose InquireMobile, a novel model inspired by reinforcement learning, featuring a two-stage training strategy and an interactive pre-action reasoning mechanism. Finally, our model achieves an 46.8% improvement in inquiry success rate and the best overall success rate among existing baselines on InquireBench. The project page is available at https://bit-aqh.github.io/InquireMobile/homepage/.
InquireMobile: Teaching VLM-based Mobile Agent to Request Human Assistance via Reinforcement Fine-Tuning
Qihang Ai1thanks: Equal contribution, random order.thanks: This work was done during an internship at Alibaba., Pi Bu [^1] footnotemark: 1, Yue Cao1, Yingyao Wang1, Jihao Gu1, Jingxuan Xing3,
Zekun Zhu3, Wei Jiang3, Zhicheng Zheng3, Jun Song1,2thanks: Corresponding author., Yuning Jiang3 1 Future Living Lab, Alibaba Group 2 Alibaba Group Holding Limited 3 Taobao & Tmall Group of Alibaba qihang005@e.ntu.edu.sg, {bupi.wj, jsong.sj}@taobao.com
1 Introduction
Recent advances in Vision-Language Models (VLMs) have significantly propelled the capabilities of mobile agents, enabling them to autonomously perceive and interact with complicated real-world environments based on human instructions Hurst et al. (2024); Bai et al. (2025); Li and Huang (2025). Traditionally, these mobile agents Hong et al. (2024); Qin et al. (2025); Deng et al. (2024) adopt a fully autonomous paradigm: after receiving user instructions, the mobile agent independently interacts with its mobile environment to accomplish designated tasks.
However, this paradigm presumes “absolute trust" in the agent’s decision-making abilities—a presupposition that becomes problematic when model understanding or reasoning capacity is insufficient. In such cases, unverified autonomy can result in severe consequences, especially in high-stakes scenarios such as online payments, order placements, or the handling of sensitive personal data as shown in Figure 1. For instance, a misinterpreted instruction at a payment interface could lead to unintended financial transactions.
@@ 1 Introduction
To bridge this research gap, we introduce InquireBench, a novel benchmark specifically designed to evaluate a mobile agent’s ability to inquire and interact safely with users. Specifically, we systematically construct diverse and challenging scenarios by leveraging both existing instructions and a vast new set of instructions generated by GPT-4o. By simulating “random walks” on real mobile phones, we trigger a breadth of potential inquire cases across 5 categories (i.e., intent confirmation, privacy and security, risk scenarios, combination and others), for evaluating user-agent interactive performance. Surprisingly, comprehensive evaluation of both open-source and closed-source VLM-based agents Bai et al. (2025); Luo et al. (2025); Lu et al. (2025b) on InquireBench reveals a near-zero baseline performance, underscoring the urgent need for new techniques tailored for user-agent collaborative situations.
To achieve this goal, we propose InquireMobile, a model designed to teach VLM-based mobile agents to request human assistance through reinforcement fine-tuning. This new mobile agent model employs a two-stage training strategy: it begins with supervised fine-tuning (SFT) for robust format acquisition, and is followed by Group Relative Policy Optimization (GRPO) Shao et al. (2024) training to enhance the model’s reasoning and thinking capabilities. Crucially, InquireMobile incorporates an interactive pre-action reasoning mechanism, where the agent proactively inquires from the user before executing critical actions—enabling informed decision-making through human-in-the-loop interaction. The pre-action mechanism is formulated as a two-step process: (i) identifying whether the current observation requires an inquiry, and (ii) generating a structured textual query to solicit user input or express uncertainty. In our experiments, InquireMobile achieves a significant performance gain of 46.8% points in inquiry success rate and the best success rate over existing baselines on InquireBench, demonstrating the feasibility and necessity of proactive user engagement in agent-driven automation.
@@
Name Eval Mode # Tasks # Apps Language Inquire Data Online
AITW Rawles et al. (2023) static - - EN
AndroidControl Li et al. (2024) static - - EN
AMEX Chai et al. (2024) static - - EN
A3 Chai et al. (2025) dynamic 201 21 EN
AppAgent Zhang et al. (2025a) dynamic 45 9 EN
AndroidArena Xing et al. (2024) dynamic 221 14 EN
AndroidWorld Rawles et al. (2024) dynamic 116 20 EN
Android-Lab Xu et al. (2024) dynamic 138 9 EN
Mobile-Env Zhang et al. (2023) dynamic 224 15 EN
CAGUI Zhang et al. (2025b) static 603 - CN
InquireBench (Ours) dynamic 173 37 EN&CN
@@ 2.1 VLM-based GUI Agents
Recent studies have shifted from complex framework designs Liu et al. (2025b) to exploring end-to-end agent strategies based on VLMs, enhancing the perception and planning capabilities of VLMs. For example, CogAgent Hong et al. (2024) presents a vision-language foundation model specializing in GUI understanding and planning, relying solely on visual inputs. UI-TARS Qin et al. (2025) continually enhances Qwen-2-VL Wang et al. (2024) in perception, reasoning, and memory by training on approximately 50 billion tokens. OS-Kairos Cheng et al. (2025) employs a confidence-driven strategy to seek help from advanced models or humans when actions are unreliable, ensuring smooth task progress. Recent benchmark studies such as AndroidLens Cao et al. (2025) further reveal that VLM-based GUI agents still struggle with long-horizon planning, memory, and robustness.
@@ 2.2 RFT-tuned GUI Agents
Inspired by the reinforcement fine-tuning (RFT) approach of DeepSeek-R1 Guo et al. (2025), RFT-tuned VLMs Zhou et al. (2025a); Liu et al. (2025c); Chen et al. (2025) have demonstrated excellent performance across various vision tasks. UI-R1 Lu et al. (2025c) and GUI-R1 Luo et al. (2025) extend this strategy to GUI agents, employing simple rule-based reward functions to evaluate the correctness of actions, achieving notable performance improvements with only limited datasets. InfiGUI-R1 Liu et al. (2025a) proposes a reasoning-oriented, two-stage training paradigm, designed to gradually evolve the agent from a reactive actor to a deliberate planner. GUI-G1 Zhou et al. (2025b) enhances the GUI interaction ability of VLMs through unified action-space rule modeling. Mobile-R1 Gu et al. (2025) further introduces an interactive multi-round reinforcement learning framework with task-level rewards, significantly improving the agent’s exploration and error-correction abilities.
Although these models have achieved impressive performance on existing benchmarks Cheng et al. (2024); Li et al. (2024, 2025), they focus on action execution while neglecting the interactive feedback from users. This poses significant challenges to the user privacy and security of GUI agents.
@@ 3.1.1 Triggering the Inquiry Situation
Existing datasets are typically constructed under idealized conditions and fail to capture the diverse range of anomalies encountered in real-world deployment. Traditional dataset construction methods typically rely on manual processes, including task design, execution planning, and human demonstration. However, such approaches are not suitable for our benchmark, as the interactive steps in our setting involve inherent randomness and unknown factors.
Inspired by RevAct Yang et al. (2025), we adopt a novel data collection strategy that reverses the conventional workflow and enables semi-automated dataset generation. Specifically, we first collect a large number of screenshots by simulating “random walks” on real mobile phones [^1] The devices used for data collection include Huawei and Xiaomi phones, both equipped with Android systems., which better reflects the natural interactions observed in daily life. Notably, to account for the significant differences in categories and usage patterns between Chinese and English apps, users are instructed to interact with each type of app separately. Each app is then evaluated within its respective system language setting.
After the “random walks”, we obtain approximately 80,345 screenshots. Human annotators are then asked to identify the interactive screenshots that require human intervention when agents encounter these screens, such as account login or payment execution. The interactive scenarios are summarized into five categories:
  • Intent Confirmation: When the agent is unable to determine the next action (e.g., pop-up windows, advertisements, or ambiguous user instructions), it proactively seeks user guidance.
  • Privacy and Security: Actions such as login or permission granting require user confirmation.
@@ 3.1.2 Thinking Generation
Annotators assess whether each item is appropriate and reasonable for the given context, and then decide whether to retain the original content, edit any unreasonable parts, or rewrite them entirely if necessary. They evaluate the model’s “thinking” for reasonableness using the following format. Redundant or incorrect data is rewritten.
<think>The user’s instruction is to send a 200-yuan red envelope in the family WeChat group. The current screenshot shows the amount entered as 200 yuan. Because sending a red envelope is a high-risk operation, user confirmation is required. </think>
The red part represents the user instruction, the blue part represents the current state or interface, and the purple part represents the reason and category in inquiry situations.
@@ 3.1.3 Statistics
The category distribution is shown in Figure 3. More details can be found in Appendix B.3.
@@ 4 InquireMobile
Our InquireMobile adopts a two-stage training strategy: (1) format finetuning by supervised fine-tuning (SFT), and (2) inquiry enhancement via GRPO training with verifiable rewards, to improve the agent’s interactive capabilities, as shown in Figure 4. The data used in the two training stages include both inquiry-related and general GUI human-annotated data, as described in Section 3. Given a task presented with a user instruction in natural language, the mobile agent is responsible for making action decisions to complete this instruction on the phone, if necessary, calls the user for assistance or clarification.
@@ 4.2 Stage 2: Rule-Based RL Training
Subsequently, the model was trained using GRPO with verifiable rewards. The reward function consists of format and action reward, defined as:
@@
Format Reward ($R_{F}$).
Following previous work Meng et al. (2025); Gu et al. (2025); Huang et al. (2025), we introduce the format reward $R_{F}$ to encourage the model to generate structured and interpretable outputs.
@@
  • <tool_call>: The final answer is a JSON object with the function name and arguments.
@@
Action Type Reward ($R_{T}$).
$R_{T}$ assesses the the correctness of the predicted action type. In our tasks, the action space consists of nine types, including general actions such as click and swipe, as well as the inquiry action call_user. The action space is detailed in Table 7 in Appendix B.2. $R_{T}$ is 1 if the action type exactly matches the ground truth otherwise, it is 0.
@@
Action Argument Reward ($R_{A}$).
  • For text-based actions (e.g., type, call_user), we compute the BLEU Papineni et al. (2002) score between the generated text and the ground truth. The BLEU score is normalized to the range $[0,1]$.
@@ 5.1 Implementation Details
The Android Studio emulator and two physical Android mobile phones serve as interaction environments. Local monitoring scripts run on the host machine and connect to each device to manage the interaction loop.
@@
In the two training stages, we utilized 975 inquiry data to enhance interactive ability and 3,000 general GUI data to retain model’s mobile agent ability. Further more, we constructed a bilingual benchmark of 190 carefully curated instructions—95 Chinese and 95 English—each crafted to elicit uncertainty and require human-in-the-loop interaction. To mirror authentic mobile usage, each task is paired with a realistic execution environment, as detailed in Appendix B.4.
@@
Qwen2.5-VL-3B is trained in two stages on 4 H100 GPUs as the base model. In Stage 1, we perform supervised fine-tuning (SFT) with LoRA (lora_rank =8=8) for 2 epochs with a learning rate of $1.0\times 10^{-4}$. In Stage 2, GRPO is applied for 2 epochs with 4 generations per sample and a temperature of 1 for exploration. The max steps is set to 15.
@@
AppAgent Zhang et al. (2025a), Mobile-Agent-E Wang et al. (2025), Qwen2.5-VL-3B Bai et al. (2025), GUI-Owl Ye et al. (2025), GUI-R1-3B Luo et al. (2025), UI-R1-3B and UI-R1-3B-E Lu et al. (2025b) are baselines.
@@
We evaluate the performance using the following metrics: 1) Task Success Rate (SR): the proportion of tasks the agent completes successfully. 2) Inquiry Success Rate (ISR): the proportion of cases in which the agent makes a correct inquiry to the user at the appropriate time. 3) Task Completion Score (Score): the score of task completion rated by GPT-4o [^2] The version we used is GPT-4o-0806. Prompts provided to GPT-4o can be found in the Appendix B.1..